OpenStudy (anonymous):

Write the equation of a vertical line that passes through the point (3,-3)

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

Write y = 1/8x + 7 in the standard form using integers.

OpenStudy (anonymous):

give medal for previous one?

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

y=(1/8)x+7 I will multiply the entire equation by 8 to get rid of the fraction.... 8y=x+56

OpenStudy (anonymous):

1/8x-y=7 x-8y=-56